Please adhere to the rounding instructions(in red) for calculations and final answer. If the answer is not exact, it will be incorrect. Thanks!

Question 11 3 pts Your firm has been working on an advanced technology. This technology will be available in the near term. The firm anticipates the first annual cash flow from the technology to be $141504 received three years from today. Subsequent annual cash flows will grow at 2.33% in perpetuity. what is the present value of the technology if the discount rate is 8.35%? (Round answer to 2 decimal places. Do not round intermediate calculations). Topic: Discounted Cash Flow Valuation
Answer #1

We have to calculate the Present Value(PV) of the technology at a discounting rate of 8.35%.

Present Value of Cash Flows that would originate form the technology = PV of Technology.

First Cash Flow is going to happen 3 yrs from now.

FV (after 3 yrs) = 141504.

Formula for Present Value of Future Cash Flow (PV) = FV / (1+r)^n

Where FV = Future Cash Flow, r = rate of Discount (Interest Rate), n= no of periods.

Present Value of $ 141504 that we would receive after 3 yrs = 141504 / (1+0.0835)^3

= $ 111245.376.

From the 4th Year we would receive $141504 every year in perpetuity with an annual growth of 2.33%.

Formula for Present Value of Growing perpetuity = FV / r-g

Where FV = Future Cash Flow, r = rate of Discount (Interest Rate), g= growth rate.

= $ 1694658.659

But the perpetual cash flow value calculated above would be after 3 yrs

Present value of the same would be 1694658.659 / (1+0.0835)^3

= $1332279.938.

Present Value of the Technology = $1443525.31 (rounded to two decimal points as required)